OpenStudy (anonymous):

Help me work this out Please I need to know the formula for this question and the work shown step by step ... Simpson drives his car with an average velocity of 48km/h to the east. How long will it take him to drive 144km on a straight highway ?

OpenStudy (anonymous):

v=48 km/h d=144 km v=d/t or t=d/v t=144/48=3 hours

OpenStudy (noelgreco):

What has Manita learned?

OpenStudy (mayankdevnani):

\[\frac{144km}{48km/hr} = 3hrs\]

OpenStudy (mayankdevnani):

@Manita